# MultiSensor-Home: A Wide-area Multi-modal Multi-view Dataset for Action Recognition and Transformer-based Sensor Fusion

### v1

@@ Abstract
Multi-modal multi-view action recognition is a rapidly growing field in computer vision, offering significant potential for applications in surveillance. However, current datasets often fail to address real-world challenges such as wide-area environmental conditions, asynchronous data streams, and the lack of frame-level annotations. Furthermore, existing methods face difficulties in effectively modeling inter-view relationships and enhancing spatial feature learning. In this study, we propose the Multi-modal Multi-view Transformer-based Sensor Fusion (MultiTSF) method and introduce the MultiSensor-Home dataset, a novel benchmark designed for comprehensive action recognition in home environments. The MultiSensor-Home dataset features untrimmed videos captured by distributed sensors, providing high-resolution RGB and audio data along with detailed multi-view frame-level action labels. The proposed MultiTSF method leverages a Transformer-based fusion mechanism to dynamically model inter-view relationships. Furthermore, the method also integrates a external human detection module to enhance spatial feature learning. Experiments on MultiSensor-Home and MM-Office datasets demonstrate the superiority of MultiTSF over the state-of-the-art methods. The quantitative and qualitative results highlight the effectiveness of the proposed method in advancing real-world multi-modal multi-view action recognition.
@@ I Introduction
Action recognition is a critical area of research in computer vision, with applications spanning surveillance [10], robotics [29], and video content analysis [20]. Traditional single-view action recognition approaches [8, 27, 13] are constrained by their reliance on a single field-of-view, resulting in incomplete contextual understanding and misclassification, particularly in cases of occlusion or partial visibility of actions. These limitations have driven the adoption of multi-view systems [19], which allow actions to be observed from multiple perspectives, enabling a more comprehensive and accurate understanding of human activities.
Multi-view action recognition integrates visual data from multiple cameras to exploit complementary perspectives and capture the full spatial context of an action. While most existing studies [33, 2, 23, 30] focus on sensor setups in narrow area coverage (Figure 1(a)), they often fail to generalize to complex, real-world scenarios where actions occur over larger areas and across viewpoints (Figure 1(b)). In wide area coverage settings, the spatial dispersion of sensors introduces additional challenges, such as targets moving across multiple views and the need to maintain consistent tracking of actions across diverse perspectives. These issues are compounded by the requirement for efficient fusion of complementary multi-view and multi-modal information while minimizing redundancy.
Recent advancements in multi-modal action recognition have highlighted the benefits of integrating diverse sensory inputs, such as audio and video [17, 6]. Existing multi-modal approaches have been limited to single-view settings, and there remains a lack of methods addressing the challenges of integrating multi-modal data in multi-view configurations. On the other hand, existing multi-modal multi-view datasets designed for wide-area sensor settings are also limited by several factors. Most notably, these datasets often provide only weak video sequence-level labels, which limits their usability for tasks requiring fine-grained spatial and temporal analysis. For instance, the MM-Office [35] and MM-Store [34] datasets lack detailed frame-level annotations, making them unsuitable for strongly supervised learning approaches. Moreover, these datasets are confined to controlled environments, lacking the diversity needed to represent real-world scenarios involving temporal dynamics and environmental variations.
@@
TABLE I: Comparison of the MM-Home dataset with existing multi-modal multi-view action recognition datasets. The MultiSensor-Home dataset offers high-resolution data with frame-level annotations.
Settings Dataset Name Year Modality #Views #Videos Avg. Duration #Classes Resolution Annotation
@@
NTU RGB+D [24] 2016 RGB+D 3 56,880   5 seconds   60 1920 $\times$ 1080 Video-level
NTU RGB+D 120 [14] 2019 RGB+D 3 114,480   5 seconds 120 1920 $\times$ 1080 Video-level
@@
Wide Area MM-Office [35] 2022 RGB+Audio 4 1,760 60 seconds   12 2560 $\times$ 1440 Video-level
MM-Store [34] 2024 RGB+Audio 6 2,970 60 seconds   18 3840 $\times$ 2160 Video-level
MultiSensor-Home (Ours) 2025 RGB+Audio 5 2,555 80 seconds   16 4000 $\times$ 3000 Frame-level
To address these challenges, we propose the Multi-modal Multi-view Transformer-based Sensor Fusion (MultiTSF) method, which is applicable to both narrow area and wide area coverage settings. MultiTSF utilizes a Transformer-based attention mechanism to dynamically model inter-view relationships and capture temporal dependencies across multiple sensors. Additionally, to address wide area coverage scenarios where targets move across dispersed views, we incorporate a Human Detection Module to enhance spatial feature learning. This module enables the model to prioritize frames and views with human activity, which is crucial for reducing redundancy and concentrating on actionable data. We also introduces the new benchmark MultiSensor-Home dataset, which provides untrimmed videos that include multiple actions captured across multiple views with detailed frame-level annotations. These videos are recorded using distributed sensors covering a wide area, including audio and RGB modalities. The proposed dataset also encompasses diverse scenarios, such as variations in time of day, clothing, and environmental conditions, making it a robust resource for studying real-world action recognition.
@@
  • MultiSensor-Home Dataset. We introduce a multi-modal multi-view dataset with fine-grained multi-view frame-level labels. This dataset addresses the limitations of existing datasets by incorporating varied environmental conditions, diverse action scenarios, and synchronized audio-visual data.
  • MultiTSF Method. We propose a method for multi-modal multi-view action recognition that combines audio and visual inputs using a Transformer-based sensor fusion mechanism to model inter-view relationships dynamically. To support spatial feature learning in MultiTSF, we introduce a Human Detection Module that generates pseudo-ground-truth annotations for human presence. This guide the model to prioritize actionable frames, which are frames containing human activity relevant to action recognition.
  • Extensive Evaluation. We evaluate MultiTSF on the proposed MultiSensor-Home dataset and MM-Office dataset [35], showing significant improvements over existing state-of-the-art methods. The quantitative and qualitative results demonstrate the effectiveness of the proposed method.
The remainder of this paper is structured as follows: Section II reviews related work. Section III and Section IV introduces the proposed MultiSensor-Home dataset and MultiTSF method, respectively. Section V details the experimental results and analysis. Finally, Section VI concludes the paper and outlines future research directions.
@@ II Related Work
Figure 2: Room layout illustrating the placement of multi-view cameras in MultiSensor-Home dataset.
@@ II-A Multi-modal Multi-view Datasets
The advancement of action recognition techniques is enabled by the development of multi-modal multi-view datasets. These datasets provide diverse perspectives and complementary modalities, allowing for a more comprehensive understanding of human actions. One of the earliest publicly available datasets, the NorthWestern-UCLA Multi-view Action 3D Dataset (NW-UCLA) [32] utilizes narrow-area object-centered sensors to record RGB+D data (RGB color and depth information) from three camera views. Following that, the NTU RGB+D [24] and its extension NTU RGB+D 120 [14] represent significant advancements by providing large-scale multi-modal multi-view recordings. These datasets offer three synchronized views and cover up to 120 action classes, making them among the most comprehensive resources for action recognition. However, these datasets focus on controlled, narrow-area environments, with trimmed videos of short duration, which limits their applicability to real-world, temporal action recognition tasks. To address the limitations of controlled settings, the TOYOTA Smarthome [4] introduced real-world home environments with RGB+D recordings captured from seven views in the dining room, kitchen, and living room. However, the dataset lacks synchronized views, as each action is clipped per view, limiting its focus to a narrow-area scene.
Recently, Yasuda et al. introduced the MM-Office [35] and MM-Store [34] datasets, which feature RGB+Audio recordings captured in office and convenience store environments, respectively. These datasets utilize distributed sensors to capture a wide-area field-of-view, where cameras partially overlap each other. This setup reflects real-world scenarios, enabling the comprehensive study of human actions across a wide range of activities. While MM-Office and MM-Store represent significant milestones in wide-area multi-view action recognition, these datasets offer only video sequence-level labels or limited frame-level annotations confined to the test set, and they are limited to weakly supervised learning tasks. To address these gaps, we propose the MultiSensor-Home dataset, a wide-area multi-modal multi-view dataset featuring untrimmed recordings captured across varying times of day (i.e., morning, afternoon, and evening). MultiSensor-Home also provides detailed multi-view frame-level labels, enabling its use in tasks such as strongly supervised action recognition at the frame level and temporal action recognition. Table I provides a comparison of the MultiSensor-Home dataset with existing multi-modal multi-view action recognition datasets.
@@ II-B Multi-modal Multi-view Action Recognition
To enable wide-range action recognition in real-world environments, recent works have introduced distributed sensor systems and advanced fusion strategies. Yasuda et al. proposed MultiTrans [35], a method designed to integrate data from distributed sensors by modeling inter-sensor relationships. However, this approach does not incorporate temporal dynamics, which are crucial for capturing the sequential nature of actions in action recognition tasks. Similarly, an extended version of MultiTrans with Guided-MELD [34] addresses the challenges of fragmented sensor observations by distilling redundant information and supplementing missing sensor data to create comprehensive event representations. On the other hand, John et al. [9] introduced a weakly supervised latent embedding model that uses view-specific latent embeddings for downstream frame-level action recognition and detection tasks. Nguyen et al. proposed MultiASL [18], which address the lack of frame-level labels by introducing an Action Selection Learning (ASL) mechanism. This ASL mechanism leverages video sequence-level annotations to generate pseudo-frame-level labels for training the network.
Although recent methods have achieved significant accuracy in multi-view action recognition, their sensor fusion strategies often do not adequately address view invariance or incorporate temporal dynamics effectively. To improve the accuracy, we propose a method named MultiTSF, which introduces a Transformer-based sensor fusion mechanism. Unlike MultiTrans [35], which focuses primarily on inter-sensor relationships, MultiTSF models the importance of sensor-level data at the frame-level by incorporating spatiotemporal features, ensuring a more detailed and context-aware representation of action dynamics. Additionally, we introduce a human detection module to enhance spatial feature learning to generate pseudo-ground-truth annotations for frames containing human presence, thereby improving the model’s ability to effectively capture relevant spatial and temporal features.
@@ III MultiSensor-Home Dataset
We introduce the MultiSensor-Home dataset [^1] The dataset will be made publicly available in the near future., a comprehensive benchmark for realistic multi-modal multi-view action recognition in indoor home environments. The dataset features untrimmed videos captured using five synchronized cameras strategically placed across a wide-area setting, as illustrated in Figure 3. This multi-view configuration enables recording actions from diverse perspectives, ensuring comprehensive coverage of spatial dynamics within the environment. Each video includes RGB and audio modalities recorded at high resolution and frame rate, offering rich multi-modal data for advanced recognition tasks.
MultiSensor-Home captures actions performed under diverse conditions, including different times of day, varying clothing styles, and natural variations in activity settings. The dataset provides detailed multi-view frame-level annotations, enabling fine-grained spatial and temporal analysis. Details of the dataset are shown in Supplemental Material, while Figure 3 showcases examples of actions captured from multiple views and Table II presenting the action classes.
TABLE II: Action Classes in the MultiSensor-Home Dataset. The column “#Events” indicates the number of occurrences.
Classes Description #Events
@@
IV Methodology
In this section, we present the proposed MultiTSF method for multi-modal multi-view action recognition. The problem definition and an overview of the proposed method are provided in Section IV-A, followed by detailed explanations of each key component in Sections IV-B,  IV-C, and IV-D.
Figure 4: Overview of the proposed MultiTSF method. It consists of: (1) Multi-modal Feature Extraction using Shared Audio Encoder and Shared Visual Encoder to extract discriminative features; (2) Human Detection Module to detect human presence and generate pseudo-ground-truth labels; (3) Temporal Modeling and Transformer-based Fusion to capture temporal dependencies and integrate spatiotemporal features for action recognition.
@@ IV-A1 Problem Definition
MultiTSF aims to predict action sequences (i.e., multi-label in the video) from the input multi-modal data captured by multi-view sensors. The input consists of a set of audio data $A=\{A_{1},A_{2},\dots,A_{N}\}$ and corresponding video data $V=\{V_{1},V_{2},\dots,V_{N}\}$, captured by $N$ distributed sensors. Each audio data $A_{i}\in\mathbb{R}^{T\times F}$ is represented as a spectrogram, where $i\in\{1,2,\dots,N\}$, $T$ is the number of frames, and $F$ is the number of frequency bins. Similarly, video data are represented as $V_{i}\in\mathbb{R}^{T\times D\times H\times W}$, where $T$ is the number of video frames, $D$ is the number of channels, and $H$ [pixels] and $W$ [pixels] represent the height and width of each frame, respectively. The objective is to extract meaningful audio-visual and temporal features to predict multi-label action outputs. Frame-level predictions are defined as $L_{t}\in\{0,1\}^{C}$ for each frame $t$, while sequence-level predictions are represented as $L\in\{0,1\}^{C}$, where $C$ is the total number of action classes.

### v3

@@ Abstract
Multi-modal multi-view action recognition is a rapidly growing field in computer vision, offering significant potential for applications in surveillance. However, current datasets often fail to address real-world challenges such as wide-area distributed settings, asynchronous data streams, and the lack of frame-level annotations. Furthermore, existing methods face difficulties in effectively modeling inter-view relationships and enhancing spatial feature learning. In this paper, we introduce the MultiSensor-Home dataset, a novel benchmark designed for comprehensive action recognition in home environments, and also propose the Multi-modal Multi-view Transformer-based Sensor Fusion (MultiTSF) method. The proposed MultiSensor-Home dataset features untrimmed videos captured by distributed sensors, providing high-resolution RGB and audio data along with detailed multi-view frame-level action labels. The proposed MultiTSF method leverages a Transformer-based fusion mechanism to dynamically model inter-view relationships. Furthermore, the proposed method integrates a human detection module to enhance spatial feature learning, guiding the model to prioritize frames with human activity to enhance action the recognition accuracy. Experiments on the proposed MultiSensor-Home and the existing MM-Office datasets demonstrate the superiority of MultiTSF over the state-of-the-art methods. Quantitative and qualitative results highlight the effectiveness of the proposed method in advancing real-world multi-modal multi-view action recognition.
@@ I Introduction
Action recognition is a critical area of research in computer vision, with applications spanning surveillance [10], robotics [29], and video content analysis [20]. Traditional single-view action recognition approaches [8, 27, 13] are constrained by their reliance on a single field-of-view, resulting in incomplete contextual understanding and misclassification, particularly in cases of occlusion or partial visibility of actions. These limitations have driven the adoption of multi-view systems [19], which allow actions to be observed from multiple field-of-views, enabling a more comprehensive and accurate understanding of human activities.
Multi-view action recognition integrates visual data from multiple cameras to exploit complementary perspectives and capture the full spatial context of an action. While most existing studies [33, 2, 23, 30] focus on sensor setups in narrow-area coverage (Figure 1(a)), they often fail to generalize to complex, real-world scenarios where actions occur over wider areas and across viewpoints (Figure 1(b)). In a wide-area coverage distributed setting, the spatial dispersion of sensors introduces additional challenges, such as targets moving across multiple views and the need to maintain consistent tracking of actions across diverse perspectives. These issues are compounded by the requirement for efficient fusion of complementary multi-view information while minimizing redundancy.
Recent advancements in multi-modal action recognition have highlighted the benefits of integrating diverse sensory inputs, such as audio and video [17, 6]. Existing multi-modal approaches have been limited to single-view setting, and there remains a lack of methods addressing the challenges of integrating multi-modal data in multi-view configurations. On the other hand, existing multi-modal multi-view datasets designed for wide-area distributed setting are limited by several factors. Most notably, these datasets often provide only weak video sequence-level labels, which limits their usability for tasks requiring fine-grained spatial and temporal analysis. For instance, the MM-Office [35] and MM-Store [34] datasets lack detailed frame-level annotations for majority of the sequences, making them unsuitable for strongly supervised learning approaches.
@@
TABLE I: Comparison of the proposed MultiSensor-Home dataset with existing multi-modal multi-view action recognition datasets.
The proposed dataset offers high-resolution data with frame-level annotations.
Setting Dataset Name Year Modality #Views #Videos Avg. Duration #Classes Resolution Annotation
@@
NTU RGB+D [24] 2016 RGB+D 3 56,880   5 seconds   60 1,920 $\times$ 1,080 Video-level
NTU RGB+D 120 [14] 2019 RGB+D 3 114,480   5 seconds 120 1,920 $\times$ 1,080 Video-level
@@
Wide Area MM-Office [35] 2022 RGB+Audio 4 1,760 60 seconds   12 2,560 $\times$ 1,440 Video-level
MM-Store [34] 2024 RGB+Audio 6 2,970 60 seconds   18 3,840 $\times$ 2,160 Video-level
MultiSensor-Home (Proposed) 2025 RGB+Audio 5 2,555 80 seconds   16 4,000 $\times$ 3,000 Frame-level
To address these challenges, we introduce a new benchmark, the MultiSensor-Home dataset, which provides untrimmed videos that include multiple actions captured across multiple views with detailed frame-level annotations. These videos are recorded using distributed sensors covering a wide-area, including audio and RGB image modalities. The proposed dataset also encompasses diverse scenarios, such as variations in time of day, clothing, and environmental conditions, making it a robust resource for studying real-world action recognition. We also propose the Multi-modal Multi-view Transformer-based Sensor Fusion (MultiTSF) method, which is applicable to both narrow-area and wide-area distributed settings. MultiTSF utilizes a Transformer-based attention mechanism to dynamically model inter-view relationships and capture temporal dependencies across multiple sensors. Additionally, to address wide-area distributed scenarios where targets move across dispersed views, we integrate a Human Detection Module to enhance spatial feature learning. This module enables the model to prioritize frames and views with human activity, which is crucial for reducing redundancy and concentrating on actionable data.
@@
  • MultiSensor-Home Dataset. We introduce a multi-modal multi-view dataset with fine-grained multi-view frame-level action labels. This dataset addresses the limitations of existing datasets by incorporating varied environmental conditions, diverse action scenarios, and synchronized audio-visual data.
  • MultiTSF Method. We propose a method for multi-modal multi-view action recognition that combines audio and visual inputs using a Transformer-based sensor fusion mechanism to model inter-view relationships dynamically. To support spatial feature learning in MultiTSF, we introduce a Human Detection Module that generates pseudo-ground-truth annotations for human presence. This guides the model to prioritize actionable frames, which are frames containing human activity relevant to action recognition.
  • Extensive Evaluation. We evaluate MultiTSF on the proposed MultiSensor-Home dataset and the existing MM-Office dataset [35], showing significant improvements over state-of-the-art methods. Quantitative and qualitative results demonstrate the effectiveness of the proposed method.
The remainder of this paper is structured as follows: Section II reviews related work. Sections III and  IV introduce the proposed MultiSensor-Home dataset and the MultiTSF method, respectively. Section V details the experimental results and analysis. Finally, Section VI concludes the paper.
@@ II Related Work
Figure 2: Room layout illustrating the placement of multi-view sensors in the proposed MultiSensor-Home dataset.
@@ II-A Multi-modal Multi-view Datasets
The advancement of action recognition techniques is enabled by the development of multi-modal multi-view datasets. These datasets provide diverse perspectives and complementary modalities, allowing for a more comprehensive understanding of human actions. One of the earliest publicly available datasets, the NorthWestern-UCLA Multi-view Action 3D (NW-UCLA) dataset [32] utilizes narrow-area object-centered sensors to record RGB+D data (RGB color and Depth information) from three camera views. Following that, the Nanyang Technological University (NTU) RGB+D [24] and its extension NTU RGB+D 120 [14] datasets represent significant advancements by providing large-scale multi-modal multi-view recordings. They offer three synchronized views and cover up to 120 action classes, making them among the most comprehensive resources for action recognition. However, these datasets focus on controlled, narrow-area setting, with trimmed videos of short duration, which limit their applicability to real-world action recognition tasks. To address the limitations of controlled settings, the Toyota Smarthome dataset [4] introduced real-world home environments with RGB+D recordings captured from seven views in a dining room, kitchen, and living room. However, it lacks synchronized views, as each action is clipped per view, limiting its focus to a narrow-area scene.
Recently, Yasuda et al. introduced the MM-Office [35] and MM-Store [34] datasets, which feature RGB+Audio recordings captured in office and convenience store environments, respectively. These datasets utilize distributed sensors to capture a wide-area field-of-view, where camera views partially overlap each other. This setup reflects real-world scenarios, enabling the comprehensive study of human actions across a wide range of activities. While these dataset represent significant milestones in wide-area multi-view action recognition, they offer only video sequence-level labels or limited frame-level annotations confined to the test set. Thus, their use is limited to weakly supervised learning tasks.
To address these gaps, we propose the MultiSensor-Home dataset, a wide-area multi-modal multi-view dataset featuring untrimmed recordings captured across varying times of day (i.e., morning, afternoon, and evening). It also provides detailed multi-view frame-level labels, enabling its use in tasks such as strongly supervised action recognition at the frame level and temporal action recognition. Table I provides a comparison of the proposed MultiSensor-Home dataset with existing multi-modal multi-view action recognition datasets.
@@ II-B Multi-modal Multi-view Action Recognition
To enable wide-area action recognition in real-world environments, recent works have introduced distributed sensor systems and advanced fusion strategies. Yasuda et al. proposed MultiTrans [35], a method designed to integrate data from distributed sensors by modeling inter-sensor relationships. However, it does not incorporate temporal dynamics, which are crucial for capturing the sequential nature of actions in action recognition tasks. Similarly, an extended version of MultiTrans with Guided-MELD [34] addresses the challenges of fragmented sensor observations by distilling redundant information and supplementing missing sensor data to create comprehensive event representations. On the other hand, John et al. [9] introduced a weakly supervised latent embedding model that uses view-specific latent embeddings for downstream frame-level action recognition and detection tasks. Nguyen et al. proposed MultiASL [18], which addresses the lack of frame-level labels by introducing an Action Selection Learning (ASL) mechanism. This mechanism leverages video sequence-level annotations to generate pseudo-frame-level labels for training the network.
Although these methods have achieved significant accuracy in multi-modal multi-view action recognition, their sensor fusion strategies often do not adequately address view invariance or incorporate temporal dynamics effectively. To improve the accuracy, we propose MultiTSF method, which introduces a Transformer-based sensor fusion mechanism. Unlike MultiTrans, which focuses on sensor relationships at the video sequence-level, MultiTSF models the importance of sensor-level data at the frame-level by incorporating spatiotemporal features, ensuring a more detailed and context-aware representation of action dynamics. Additionally, we introduce a Human Detection Module to enhance spatial feature learning to generate pseudo-ground-truth annotations for frames containing human presence, thereby improving the ability of the model to effectively capture relevant spatial and temporal features.
@@ III MultiSensor-Home Dataset
TABLE II: Action Classes in the proposed MultiSensor-Home Dataset. “#Events” indicates the number of occurrences.
Class Description #Events
@@
Figure 4: Overview of the proposed MultiTSF method. It consists of: (1) Multi-modal Feature Extraction using Shared Audio Encoder and Shared Visual Encoder to extract discriminative features, (2) Human Detection Module to detect human presence and generate pseudo-ground-truth labels, and (3) Temporal Modeling and Transformer-based Fusion to capture temporal dependencies and integrate spatiotemporal features for action recognition.
We introduce the MultiSensor-Home dataset, a comprehensive benchmark for realistic multi-modal multi-view action recognition in indoor home environments. The dataset features untrimmed videos captured using five synchronized cameras strategically placed across a wide-area distributed setting, as illustrated in Figure 3, where a single person performs actions within the room. This multi-view configuration enables recording actions from diverse perspectives, ensuring comprehensive coverage of spatial dynamics within the environment. Each video includes RGB and audio modalities, recorded at a 4K resolution of $4,000\times 3,000$ pixels and a frame rate of 30 frames per second, offering high-quality multi-modal data for action recognition tasks. Moreover, MultiSensor-Home captures actions performed under diverse conditions, including different times of day, varying clothing styles, and natural variations in activity settings. The dataset provides detailed multi-view frame-level annotations, enabling fine-grained spatial and temporal analysis. Figure 3 showcases examples of actions captured from multiple views and Table II presenting the action classes.
IV Multi-modal Multi-view
Transformer-based Sensor Fusion
In this section, we propose the MultiTSF method for multi-modal multi-view action recognition. The problem definition and an overview of the proposed method are provided in Section IV-A, followed by detailed explanations of each key component in Sections IV-B,  IV-C, and IV-D.
@@ IV-A1 Problem Definition
MultiTSF aims to predict action sequences (i.e., multi-label in the video) from the input multi-modal data captured by multi-view sensors (views). The input consists of a set of audio data $A=\{A_{1},A_{2},\dots,A_{N}\}$ and corresponding video data $V=\{V_{1},V_{2},\dots,V_{N}\}$, captured by $N$ distributed sensors. Each audio data $A_{n}\in\mathbb{R}^{T\times F}$ is represented as a spectrogram, where $n\in\{1,2,\dots,N\}$, $T$ is the number of frames, and $F$ is the number of frequency bins. Similarly, video data are represented as $V_{n}\in\mathbb{R}^{T\times D\times H\times W}$, where $T$ is the number of video frames, $D$ is the number of channels, and $H$ [pixels] and $W$ [pixels] represent the height and width of each frame, respectively. The objective is to extract meaningful audio-visual and temporal features to predict multi-label action outputs. Frame-level predictions are defined as $L_{t}\in\{0,1\}^{C}$ for each frame $t\in\{1,2,\dots,T\}$, while sequence-level predictions are represented as $L\in\{0,1\}^{C}$, where $C$ is the total number of action classes.